AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Thu Apr 3 19:29:17 2025
Author
*
:
Subject
*
:
15.40 DSSSD bias & leakage current OK - attachments 1-2 log shows variations due to ambient temperature changes in CRYRING hall FEE64 temperatures OK - attachment 3 WR timestamps OK - attachment 4 System wide checks Ok *except* global clock (aida16) - attachments 5-6 ADC data item stats OK - attachment 7 per FEE64 Rate spectra - attachment 8 per FEE64 1.8.L spectra - attachments 9-10 have not been cleared in some time probably - will zero histograms & stats for later review per FEE64 1.8.W spectra - 20us FSR - attachments 11-12 Merger, TapeServer etc - attachments 13-14 data file G22-222/R51_15 Motion control - attachment 15 16.01 all histograms & stats zero'd 16.43 online per DSSSD energy spectra - attachment 16 photographs of CRYRING beam & target diagnostics - attachments 17-22 18.45 Left DSSSDs cycled to clear high rates (c. 250k/FEE64) aida01-aida04 per https://elog.ph.ed.ac.uk/CARME/603 18.49 DAQ continues data file G22-201/R51_27 DSSSD bias & leakage current OK - attachment 23 FEE64 temperatures OK - attachment 24 ADC data item stats OK - attachment 25 per FEE64 Rate spectra - attachment 26 online per DSSSD energy spectra - attachment 27 19.00 Beam less stable but most cycles 5e+7 or higher Target peaking c. 1e12 dropping to c. 6e11 19.09 CRYRING hall temperature increasing, DSSSD leakage current increasing - attachments 28-29 20.06 Beam less stable most cycles 1-2e+7 or higher Target c. 1-2e12 20.15 JM/CB advise changing Lakeshore 336 Temp Controller set point from 29k to 29.5K - done Average pressure decreases but rapid variations still observed - see attachment 30 21.00 contacted HKR to improve beam stability - they were alreday working on it but did not know what was causing the problem - will contact on-call CRYRING expert cycled L DSSSD to reset rates 21,10 beam 0ff - DSSDs OUT Increased target temperature to 30K to solve instability that developed again up to E12 density range.
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown